For a ballistic study, a 2.4 - gm bullet is fired into soft wood. The bullet strikes the wood surface with a velocity of 358 m/s and penetrates 0.19 m. Find the constant retarding force in N. Include the sign (+ or -) if the unknown variable is a vector quantity.


1
Expert's answer
2021-02-28T07:32:37-0500

We can find the constant retarding force from the work-kinetic energy theorem:


"\Delta KE=W,""KE_f-KE_i=Fd,""0-\dfrac{1}{2}mv_i^2=Fd,""F=\dfrac{-\dfrac{1}{2}mv_i^2}{d},""F=\dfrac{-\dfrac{1}{2}\cdot2.4\cdot10^{-3}\ kg\cdot(358\ \dfrac{m}{s})^2}{0.19\ m}=-809.46\ N."

The sign minus means that the constant retarding force acts in the opposite direction to the motion of the bullet.